OpenStudy (anonymous):

simplifying the expression (–1 + i) + (21 + 5i)?

OpenStudy (anonymous):

that's -1 + i + 21 + 5i = 20 + 6i

OpenStudy (anonymous):

(–1 + i) + (21 + 5i) + 0 –1 + (i + 21) + 5i (–1 + 21) + (i + 5i) –(1 – i) + (21 + 5i)

OpenStudy (anonymous):

which one of those?

OpenStudy (anonymous):

3rd one

OpenStudy (anonymous):

what we are doing..is putting together the real numbers, and then grouping together the imaginary numbers

OpenStudy (anonymous):

help me out Which property of multiplication is shown below? If x = a + bi and y = c + di, x × y = y × x commutative property identity property distributive property associative property